OpenStudy (anonymous):

What are the solutions of 2x^2 + 3x = -5

OpenStudy (anonymous):

\[A. -3 \pm \sqrt{31}\div4\]

OpenStudy (anonymous):

\[B. -3\pm7i \div4\]

OpenStudy (anonymous):

C. \[C. -3\pm i \sqrt{19}\div4\]

OpenStudy (anonymous):

\[D. -3\pm4i \div4\]

OpenStudy (shinalcantara):

use quadratic formula \[x = \frac{ -b \pm \sqrt{b ^{2}-4ac}}{ 2a }\]

OpenStudy (haseeb96):